OpenStudy (anonymous):

The length of the base of a square pyramid is 9 inches. The slant height is 12 inches. What is the surface area of the pyramid?

Directrix (directrix):

The surface area of a regular pyramid is (1/2) the perimeter of the base (square in this case) times time slant height plus the area of the base.

Directrix (directrix):

SA = (1/2)* (36) * 12 + 9*9 where 36 is the base perimeter and 9*9 is the area of the base. SA = (1/2)* (36) * 12 + 9*9 SA = 18 * 36 + 81 SA = ? @jawan0513